jeff and jacks ages are such that twice jeff's present age is the same as jacks ge was 4 years ago. jeff is now one third as old as jack will be 26 years from now. find the current age of both jeff and jack
Answer by lwsshak3(11628)

jeff and jacks ages are such that twice jeff's present age is the same as jacks ge was 4 years ago. jeff is now one third as old as jack will be 26 years from now. find the current age of both jeff and jack
***
let x=jeff's present age
2x+4=jack's present age
3x-26=also, jack's present age
2x+4=3x-26
x=30
2x+4=64
..
4 yrs ago:
2x=jack's age
..
26 yrs from now:
3x=jack's age
jeff's current age=30
jack's current age=64